poloololol:

Timothy recorded the monthly change in his savings account balance, in dollars, for 8 months. Month January February March April May June July August Balance change ($) −30 −15 55 40 −20 −45 −10 65 What is the average monthly change in Timothy's savings account balance?

Bones:

What's given in the problem. Timothy recorded the monthly change in his savings account balance, in dollars, for 8 months The monthly changes are -30, -15, 55,40,-20, -45, -10, and 65 How to solve Calculate the average by summing the changes and dividing by the number of months. Step 1 …Sum the monthly changes. -30+-15+55+40+-20+-45+-10+65=40 Step 2 …Divide the sum by the number of months (8). 40/8=5 Solution The average monthly change in Timothy's savings account balance is $5

Bones:

Hope this helps

MAGABACK:

@bones wrote:
What's given in the problem. Timothy recorded the monthly change in his savings account balance, in dollars, for 8 months The monthly changes are -30, -15, 55,40,-20, -45, -10, and 65 How to solve Calculate the average by summing the changes and dividing by the number of months. Step 1 …Sum the monthly changes. -30+-15+55+40+-20+-45+-10+65=40 Step 2 …Divide the sum by the number of months (8). 40/8=5 Solution The average monthly change in Timothy's savings account balance is $5
dang.